### v2

@@ 2.3 Framework structure
(b) The surrogate model of $\boldsymbol{\alpha}_{3}(\mathbf{X}_{j},\mathbf{X}_{k},\mathbf{X}_{i})$ is the MLP neural network $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$ obtained through 3-CliqueConv (referred to as FaceConv in the schematic following [30]).
@@ 2.5 Prediction: transferability and cost
HIGNN can potentially achieve significant computational speedup by requiring only a forward pass through its MLP neural networks at each time step, in contrast to traditional simulation methodologies [32, 15, 16, 18, 11, 24], which usually require solving a large-scale linear system at each time step. The computational complexity of HIGNN is directly related to the number of edges and 3-cliques in the graph, as neural network inferences are performed on each edge and 3-clique. As discussed above, the two-body HI not only dominates the short-range lubrication effect but also decays very slowly ($\sim O(r^{-1})$) in long range. Thus, for each particle, we need to consider all other particles as its neighbors for the calculation of $\boldsymbol{\alpha}^{(t)}_{2}$. In HIGNN, this translates to the construction of edges between any two particles, leading to a fully-connected graph characterized by N(N1)N(N-1) edges. As for the 3-cliques, which characterize three-body connections, the faster decaying rate of three-body HIs allows for the implementation of a finite cutoff distance ($R_{\text{cut}}$), thereby limiting the number of 3-cliques to $O(N)$. Therefore, the overall computational complexity (prediction cost) of HIGNN scales as $O(N^{2})$ with $N$ the total number of particles, which is greater than the scaling $O(N\log N)$ of ASD [24]. This implies that while HIGNN’s computational cost is significantly lower than ASD for small- to moderate-scale problems (as shown in our previous work [30, 31]), for simulations of very large-scale particulate suspensions, HIGNN’s cost can approach or surpass that of ASD. Therefore, to achieve a consistent superiority on computational efficiency, in the present work we aim to reduce the scaling of the prediction cost of HIGNN down to quasi-linear, i.e., ${O}(N\log N)$. By such, even for simulating very large-scale particulate suspensions (e.g., containing more than millions of particles), the computational cost can still be maintained low. For achieving that, we propose to integrate the $\mathcal{H}$-matrix techniques [33, 34] into the HIGNN framework. The detailed methodology is introduced in the next section.
@@ 3 Hierarchical Matrix ($\mathcal{H}$-Matrix)-Accelerated HIGNN
where $\mathbf{M}^{(1)}$, $\mathbf{M}^{(s,2)}$ and $\mathbf{M}^{(s,3)}$ are all diagonal matrices with each element (a $3\times 3$ submatrix) calculated from $\boldsymbol{\alpha}_{1}$, $\boldsymbol{\alpha}^{(s)}_{2}$ and $\boldsymbol{\alpha}^{(s)}_{3}$, respectively; the matrix $\mathbf{M}^{(t,2)}$ is constructed from the two-body interactions $\boldsymbol{\alpha}_{2}^{(t)}$, and $\mathbf{M}^{(t,3)}$ from the three-body interactions $\boldsymbol{\alpha}_{3}^{(t)}$. Following the preceding analysis, the matrix-vector multiplications appearing in Eq. (6) typically exhibit $O(N)$ complexity due to the sparsity of the associated matrices, but with the exception of the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$, which has $O(N^{2})$ complexity because $\mathbf{M}^{(t,2)}$ is a dense matrix.
We determined that neither the Particle-Mesh Ewald (PME) method nor the Fast Multipole Method (FMM) were suitable for our problem. The PME method was originally designed for periodic boundary conditions [45] and achieves optimal performance in this context. While extensions were developed for non-periodic and semi-periodic systems [46, 47], these modifications could introduce additional computational overhead and errors. Thus, PME is not suitable for our problem scope, which aims to develop a method applicable to any type of boundary condition. Moreover, applying PME in our work would necessitate computing Fourier transforms of neural network-based interaction terms, introducing substantial additional computational expense. FMM typically requires analytical expansions, such as multipole, Legendre, or Taylor series [48, 49]. However, $\boldsymbol{\alpha}_{2}^{(t)}(\mathbf{X}_{i},\mathbf{X}_{j})$ is modeled by a neural network herein, and hence it is infeasible to represent $\boldsymbol{\alpha}_{2}^{(t)}$ using any analytical expansions. While it is possible to approximate $\boldsymbol{\alpha}_{2}^{(t)}$ with a sum of functions that have known multipole or other analytical expansions, this approach may compromise accuracy and incur additional computational cost due to the need to sample data from the neural network to construct the approximation. Variants like the kernel-independent FMM (KIFMM) [50] replace analytical expansions with interpolation (e.g., Chebyshev or polynomial), but its interpolation overhead can be also significant for neural network kernel evaluations. Finally, while FMM generally requires less setup time, $\mathcal{H}$-matrix tends to offer superior efficiency for matrix-vector multiplications and better parallel scalability [49, 51, 52], making it more suitable for the large-scale systems and numerous matrix-vector operations central to this work.
Thus, we integrate the $\mathcal{H}$-matrix techniques [33, 34] to reduce the computational complexity for the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$ from $O(N^{2})$ to $O(N\log N)$. Note that $\mathbf{M}^{(t,2)}$ is in fact assembled by $N\times N$ submatrices, where each submatrix $\mathbf{M}^{(t,2)}_{ij}\in\mathbb{R}^{3\times 3}$ corresponds to $\boldsymbol{\alpha}_{2}^{(t)}(\mathbf{X}_{i},\mathbf{X}_{j})$. Following the convention of $\mathcal{H}$-matrix, it is more convenient to let the matrix’s dimension equal to the number of particles. Hence, we denote each submatrix $\mathbf{M}^{(t,2)}_{ij}$ as an element or entry of the matrix, and thereby $\mathbf{M}^{(t,2)}$ is referred to as an $N\times N$ matrix. [^1] For the subsequent discussion, it should be noted that when performing any operation on elements, each element corresponds to a $3\times 3$ submatrix, rather than a scalar value.
@@ 3.1 Cluster-tree block partitioning
The process begins at the root of the cluster tree, where $\boldsymbol{\tau}=\boldsymbol{\sigma}=\mathbf{I}$. For each candidate block $(\boldsymbol{\tau},\boldsymbol{\sigma})$, the admissibility condition is first evaluated. If the condition is met, the block is added to the set of admissible blocks, $\mathcal{P}^{G}$. Otherwise, the algorithm recursively examines all sons of both $\boldsymbol{\tau}$ and $\boldsymbol{\sigma}$ until all leaves are reached, i.e. $|\boldsymbol{\tau}|<C_{\text{leaf}}$ and $|\boldsymbol{\sigma}|<C_{\text{leaf}}$. Any leafs that do not satisfy the admissibility condition are added into the set of inadmissible blocks, $\mathcal{P}^{H}$. Finally, the complete sets of $\mathcal{P}^{G}$ and $\mathcal{P}^{H}$ are obtained. Figure 4 illustrates this process using a one-dimensional example with the root $\mathbf{I}=[0,1]$ and the admissibility parameter $\eta=1$.
@@ 4.2 Assessment of complexity and scalability
We next assess the computational complexity of the $\mathcal{H}$-HIGNN model. It can be evaluated by examining the total number of submatrices inferred through GNN during the construction of the $\mathcal{H}$-HIGNN model. Thus, we compare the number of queries for two-body HIs $\boldsymbol{\alpha}_{2}^{(t)}$ by $\mathcal{H}$-matrix with the combined sum of other HIs, i.e. $\boldsymbol{\alpha}_{2}^{(s)}$, $\boldsymbol{\alpha}_{3}^{(t)}$ and $\boldsymbol{\alpha}_{3}^{(s)}$, as shown in Figure 7. We extended the system size to 10 million particles for this section’s analysis, enabling a more comprehensive investigation on the complexity and scalability. The results in Figure 7 reveal that $\boldsymbol{\alpha}_{2}^{(t)}$ dominates the overall computational complexity, exhibiting 1-2 orders of magnitude more queries than the other interaction terms. And we are able to reduce its complexity down to $O(N\log N)$ from $O(N^{2})$ through integrating $\mathcal{H}$-matrix. In light of its dominant role, in the following tests, we focus on the complexity and scalability of $\boldsymbol{\alpha}_{2}^{(t)}$ under varying $\mathcal{H}$-matrix parameters.
@@
Figures 8 and 9 present the computational complexity analysis across different parameter settings. Using the same two scenarios for particle configurations as described in §4.1, the results demonstrate that the $\mathcal{H}$-matrix consistently achieves a computational complexity of $O(N\log N)$, independent of parameter choices. This represents a substantial improvement over the original HIGNN model’s quadratic scaling of $O(N^{2})$, leading to significant enhancement in computational efficiency, particularly for large-scale systems. For systems containing 10 million ($10^{7}$) particles, integrating $\mathcal{H}$-matrix into HIGNN reduces the number of required queries by 3-4 orders of magnitude. Furthermore, when considering these results in together with the error analysis, improving the accuracy of $\mathcal{H}$-matrix is more effectively achieved by reducing $\zeta$ rather than increasing $C_{\text{leaf}}$. A decrease in $\zeta$ yields better accuracy improvements while incurring a smaller computational overhead. Therefore, in practical applications, one can optimize performance by selecting smaller leaf node limits to enhance computational efficiency while utilizing $\zeta$ as the primary parameter for error control.

### v3

@@ 2.3 Framework structure
(b) The surrogate model of $\boldsymbol{\alpha}_{3}(\mathbf{X}_{j},\mathbf{X}_{k},\mathbf{X}_{i})$ is the MLP neural network $\mathbf{g}_{\boldsymbol{\Theta}_{3}}$ obtained through 3-CliqueConv.
@@ 2.5 Prediction: transferability and cost
HIGNN can potentially achieve significant computational speedup by requiring only a forward pass through its MLP neural networks at each time step, in contrast to traditional simulation methodologies [32, 15, 16, 18, 11, 24], which usually require solving a large-scale linear system at each time step. The computational complexity of HIGNN is directly related to the number of edges and 3-cliques in the graph, as neural network inferences are performed on each edge and 3-clique. As discussed above, the two-body HI not only dominates the short-range lubrication effect but also decays very slowly ($\sim O(r^{-1})$) in long range. Thus, for each particle, we need to consider all other particles as its neighbors for the calculation of $\boldsymbol{\alpha}^{(t)}_{2}$. In HIGNN, this translates to the construction of edges between any two particles, leading to a fully-connected graph characterized by N(N1)N(N-1) edges. As for the 3-cliques, which characterize three-body connections, the faster decaying rate of three-body HIs allows for the implementation of a finite cutoff distance ($R_{\text{cut}}$), thereby limiting the number of 3-cliques to $O(N)$. Therefore, the overall computational complexity (prediction cost) of HIGNN scales as $O(N^{2})$ with $N$ the total number of particles, which is greater than the scaling $O(N\log N)$ of ASD [24]. This implies that while HIGNN’s computational cost is significantly lower than ASD for small- to moderate-scale problems (as shown in our previous work [30, 31]), for simulations of very large-scale particulate suspensions, HIGNN’s cost can approach or surpass that of ASD. Therefore, to achieve a consistent superiority on computational efficiency, in the present work we aim to reduce the scaling of the prediction cost of HIGNN to quasi-linear, i.e., ${O}(N\log N)$. By such, even for simulating very large-scale particulate suspensions (e.g., containing more than millions of particles), the computational cost can still be maintained low. For achieving that, we propose to integrate the $\mathcal{H}$-matrix techniques [33, 34] into the HIGNN framework. The detailed methodology is introduced in the next section.
@@ 3 Hierarchical Matrix ($\mathcal{H}$-Matrix)-Accelerated HIGNN
where $\mathbf{M}^{(1)}$, $\mathbf{M}^{(s,2)}$ and $\mathbf{M}^{(s,3)}$ are all diagonal matrices with each element (a $3\times 3$ submatrix) calculated from $\boldsymbol{\alpha}_{1}$, $\boldsymbol{\alpha}^{(s)}_{2}$ and $\boldsymbol{\alpha}^{(s)}_{3}$, respectively; the matrix $\mathbf{M}^{(t,2)}$ is constructed from the two-body interactions $\boldsymbol{\alpha}_{2}^{(t)}$, and $\mathbf{M}^{(t,3)}$ from the three-body interactions $\boldsymbol{\alpha}_{3}^{(t)}$. Following the preceding analysis, the matrix-vector multiplications appearing in Eq. (6) typically exhibit $O(N)$ complexity due to the sparsity of the associated matrices, but with the exception of the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$, which has $O(N^{2})$ complexity because $\mathbf{M}^{(t,2)}$ is a dense matrix. Thus, we integrate the $\mathcal{H}$-matrix techniques [33, 34] to reduce the computational complexity for the multiplication $\mathbf{M}^{(t,2)}\cdot\mathbf{F}$ from $O(N^{2})$ to $O(N\log N)$. Note that $\mathbf{M}^{(t,2)}$ is in fact assembled by $N\times N$ submatrices, where each submatrix $\mathbf{M}^{(t,2)}_{ij}\in\mathbb{R}^{3\times 3}$ corresponds to $\boldsymbol{\alpha}_{2}^{(t)}(\mathbf{X}_{i},\mathbf{X}_{j})$. Following the convention of $\mathcal{H}$-matrix, it is more convenient to let the matrix’s dimension equal to the number of particles. Hence, we denote each submatrix $\mathbf{M}^{(t,2)}_{ij}$ as an element or entry of the matrix, and thereby $\mathbf{M}^{(t,2)}$ is referred to as an $N\times N$ matrix. [^1] For the subsequent discussion, it should be noted that when performing any operation on elements, each element corresponds to a $3\times 3$ submatrix, rather than a scalar value.
@@ 3.1 Cluster-tree block partitioning
The process begins at the root of the cluster tree, where $\boldsymbol{\tau}=\boldsymbol{\sigma}=\mathbf{I}$. For each candidate block $(\boldsymbol{\tau},\boldsymbol{\sigma})$, the admissibility condition is first evaluated. If the condition is met, the block is added to the set of admissible blocks, $\mathcal{P}^{G}$. Otherwise, the algorithm recursively examines all sons of both $\boldsymbol{\tau}$ and $\boldsymbol{\sigma}$ until all leaves are reached, i.e. $|\boldsymbol{\tau}|<C_{\text{leaf}}$ and $|\boldsymbol{\sigma}|<C_{\text{leaf}}$. Any leaves that do not satisfy the admissibility condition are added into the set of inadmissible blocks, $\mathcal{P}^{H}$. Finally, the complete sets of $\mathcal{P}^{G}$ and $\mathcal{P}^{H}$ are obtained. Figure 4 illustrates this process using a one-dimensional example with the root $\mathbf{I}=[0,1]$ and the admissibility parameter $\eta=1$.
@@ 4.2 Assessment of complexity and scalability
We next assess the computational complexity of the $\mathcal{H}$-HIGNN model. It can be evaluated by examining the total number of submatrices inferred through GNN during the construction of the $\mathcal{H}$-HIGNN model. Thus, we compare the number of queries for two-body HIs $\boldsymbol{\alpha}_{2}^{(t)}$ by $\mathcal{H}$-matrix with the combined sum of other HIs, i.e. $\boldsymbol{\alpha}_{2}^{(s)}$, $\boldsymbol{\alpha}_{3}^{(t)}$ and $\boldsymbol{\alpha}_{3}^{(s)}$, as shown in Figure 7. We extended the system size to 10 million particles for this section’s analysis, enabling a more comprehensive investigation on the complexity and scalability. The results in Figure 7 reveal that $\boldsymbol{\alpha}_{2}^{(t)}$ dominates the overall computational complexity, exhibiting 1-2 orders of magnitude more queries than the other interaction terms. And we are able to reduce its complexity to $O(N\log N)$ from $O(N^{2})$ through integrating $\mathcal{H}$-matrix. In light of its dominant role, in the following tests, we focus on the complexity and scalability of $\boldsymbol{\alpha}_{2}^{(t)}$ under varying $\mathcal{H}$-matrix parameters.
@@
Figures 8 and 9 present the computational complexity analysis across different parameter settings. Using the same two scenarios for particle configurations as described in §4.1, the results demonstrate that the $\mathcal{H}$-matrix consistently achieves a computational complexity of $O(N\log N)$, independent of parameter choices. This represents a substantial improvement over the original HIGNN model’s quadratic scaling of $O(N^{2})$, leading to significant enhancement in computational efficiency, particularly for large-scale systems. For systems containing 10 million ($10^{7}$) particles, integrating $\mathcal{H}$-matrix into HIGNN reduces the number of required queries by 3-4 orders of magnitude. Furthermore, when considering these results together with the error analysis, improving the accuracy of $\mathcal{H}$-matrix is more effectively achieved by reducing $\zeta$ rather than increasing $C_{\text{leaf}}$. A decrease in $\zeta$ yields better accuracy improvements while incurring a smaller computational overhead. Therefore, in practical applications, one can optimize performance by selecting smaller leaf node limits to enhance computational efficiency while utilizing $\zeta$ as the primary parameter for error control.
@@ 5 Conclusion
We note that the particle–mesh Ewald (PME) method [57, 58, 59] and the fast multipole method (FMM) [60, 61, 62] may potentially be adapted to address our problem. However, we adopted the $\mathcal{H}$-matrix approach based on our expertise. Moreover, $\mathcal{H}$-matrix tends to offer better parallel efficiency [63], which is critical for the large-scale systems considered in this work.